OpenStudy (anonymous):

Anybody good with calculus product rule? (x-2)e^(2-x)

OpenStudy (mertsj):

\[(x-2)(e ^{2-x})(-1)+e ^{2-x}(1)=(e ^{2-x})(2-x)\]

OpenStudy (mertsj):

Assuming that you want the first derivative.

OpenStudy (anonymous):

Where does the (-1) come from??

sam (.sam.):

comes from chain rule.

OpenStudy (turingtest):

what is the derivative of\[2-x\]?

OpenStudy (anonymous):

But I thought e differentiates to itself?

OpenStudy (turingtest):

but you need to use the chain rule as .Sam. said....

OpenStudy (turingtest):

\[{d\over dx}e^{2-x}=e^{2-x}\cdot{d\over dx}(2-x)=-e^{2-x}\]

OpenStudy (turingtest):

@shallster are you familiar with the chain rule? or is the post above confusing you?

OpenStudy (anonymous):

pretty confusing.

OpenStudy (turingtest):

ok, can you take the derivative of\[(2-x)^2\]?

OpenStudy (anonymous):

it's not asking me to do that though?? The equation is (x-2)e^(2-x) I need the derivatives of x-2 & e^(2-x) don't I??

OpenStudy (turingtest):

yes, but you seem to have trouble differentiating\[e^{2-x}\]so I'm trying to help you do so

OpenStudy (anonymous):

but i thought e differentiates to itself?

OpenStudy (turingtest):

again: you \(must\) use the chain rule here \[\frac d{dx}e^x=e^x\]yes, but this is not the case here, we have a compound function so we must use the chain rule

OpenStudy (anonymous):

You obviously know what you're talking about, so I'm all ears. I NEED help.

OpenStudy (turingtest):

so again I ask you to demonstrate your understanding of the chain rule what if I asked you to differentiate\[(1-x^2)^2\]by using the chain rule, could you do that?

OpenStudy (anonymous):

4x(1-x^2) ???

OpenStudy (turingtest):

exactly, you had to differentiate the inside function as well now that I know you understand this I can explain the other derivative...

OpenStudy (turingtest):

\[\frac d{dx}e^{f(x)}=e^{f(x)}\cdot f'(x)\]see? you still leave the e^(whatever) part the same, but you still must differentiate the second function, as always in the chain rule

OpenStudy (turingtest):

do you agree with the above line? let me know what is confusing you if not

OpenStudy (anonymous):

so e^(2-x) = -1e^(2-x)

OpenStudy (turingtest):

yep :) now does mertsj's answer make sense?

OpenStudy (anonymous):

I've got it down to e^2-x) - (x-2)e^(2-x) and I'm not that sure how to simplify it.

OpenStudy (turingtest):

dang lag... factor out \(e^{2-x}\) like so\[(x-2)e^{2-x}\]\[e^{2-x} -(x-2)e^{2-x}=e^{2-x}(1-x+2)=e^{2-x}(3-x)\]so I think mertsj was a little bit off

OpenStudy (anonymous):

I also struggle with factoring. I'm such a douche.

OpenStudy (turingtest):

no, this kind of factoring is a bit tricky I learned it through problems like these as you can see, mertsj made a small mistake in factoring as well, despite his illustrious abilities it happens to the best of us ;) PS: as a mod, I have to ask you not to say the "d" word above I'd like to keep you around after all!

OpenStudy (anonymous):

Understood.

OpenStudy (mertsj):

I can't believe I did that. But, shallster, as you will learn, Turing is the resident expert. So always pay attention to him.

OpenStudy (anonymous):

That I will.

OpenStudy (turingtest):

thanks y'all ! you guys make the site what it is though agape love =)

OpenStudy (mertsj):

Right back to you, Turing.

OpenStudy (anonymous):

Still not sure how the -2 became +2 when factoring?

OpenStudy (mertsj):

By distributing the -1

OpenStudy (mertsj):

\[e ^{2-x}-(x-2)e ^{2-x}=e ^{2-x}[1-(x-2)]=e ^{2-x}[1-x+2]=e ^{2-x}[3-x]\]

OpenStudy (anonymous):

I understand the rest but how does (1-(x-2)) expand to (1-x+2)?

OpenStudy (turingtest):

distribute -1

OpenStudy (anonymous):

I bet it's something really simple that my brain is bypassing!!

OpenStudy (turingtest):

like that^ ? :)

OpenStudy (turingtest):

\[-(x-2)=-x+2\]no?

OpenStudy (anonymous):

ah, i see. Thanks guys.